OpenStudy (anonymous):

3(n+2)-4(n-1)=n+13, show me how you would work this out

OpenStudy (anonymous):

n=1/2

OpenStudy (anonymous):

https://www.mathway.com/

OpenStudy (anonymous):

3n + 6 - 4n -4 = n+13

OpenStudy (anonymous):

-2n = 13 -6 +4

OpenStudy (anonymous):

I believe u can do it now

OpenStudy (anonymous):

-2n = 13 -6 +4=-5.5

OpenStudy (anonymous):

yeah

OpenStudy (texaschic101):

3(n + 2) - 4(n - 1) = n + 13 --- distribute through the parenthesis 3n + 6 - 4n + 4 = n + 13 -- combine like terms -n + 10 = n + 13 -- subtract n from both sides -n -n + 10 = 13 -- subtract 10 from both sides -n - n = 13 - 10 -2n = 3 n = -3/2 or - 1.5

OpenStudy (anonymous):

You rock texas

OpenStudy (texaschic101):

lol...thanks